OpenStudy (anonymous):

Three boys are sharing a cake. Joey ate 2/5 of the cake, Billy ate 1/3 of it, and Matt ate the remaining piece. What is the ratio of Joey's piece to Billy's piece to Matt's piece?

OpenStudy (dumbcow):

2/5 + 1/3 = 6/15 + 5/15 = 11/15 Matt piece was 4/15 of cake. getting rid of common denominator of 15 you get ratios: 6:5:4

OpenStudy (anonymous):

@dumbcow thank you soooooooooooo much!!

OpenStudy (dumbcow):

yw